OpenStudy (anonymous):

A $3,500.00 principal earns 3% interest, compounded semiannually. After 20 years, what is the balance of the account?

OpenStudy (anonymous):

Well using the interest formula A=P(1+r/n)^nt

OpenStudy (anonymous):

A=3500(1+.03/2)^(2)(20)

OpenStudy (anonymous):

=3500(1+.015)^40 =3500(1.015)^40 A=6349.06

OpenStudy (anonymous):

That should be it.